# DiffMap: Enhancing Map Segmentation with Map Prior Using Diffusion Model

### v1

@@ IV-A4 Inference Details
The prediction results are obtained by performing the denoising process $20$ times on the noisy maps under the condition of current BEV features. Considering the stochastic nature of the generative model, we perform the sampling process $3$ times, and use the mean value of sampling results as the final prediction result.
@@ IV-B2 Instance Metrics
where $C_{1}$ and $C_{2}$ are sets of points on the predicted curve and ground-truth curve. The typical Chamfer distance(CD) is bi-directional. And the CD between the predicted curve and the ground-truth curve is given by:
@@ IV-B3 Evaluation on multiple intervals
To better explore the effectiveness of DiffMap in long-range detection, we further split the ground truth into three intervals: 0-30m, 30-60m, and 60-90m to compared with SuperFusion fairly. Subsequently, within these three intervals, we calculate the IoU, CD, and AP for different methods, aiming for a comprehensive evaluation of the detection results.
Figure 4: Qualitative results: Diffmap is capable of capture the structured prior and achieves the best segmentation results. The results demonstrate effective restoration of parallel shapes for pedestrian crossings, smoothness and continuity of dividers, and shape complementation for boundaries.
@@ IV-C Evaluation Results
To facilitate better comparison, we reproduced the predominant methods for map segmentation like HDMapNet and SuperFusion algorithms on nuScenes dataset. Compared with Superfuion, Table I shows the comparisons of the IoU scores of semantic map segmentation. Our DiffMap shows a significant improvements on all intervals and has best results, especially in lane divider and pedestrian crossing. As shown in Table II, our method also demonstrates a significant improvement in terms of average precision (AP), verifying the effectiveness of our DiffMap. Besides, we incorporate our DiffMap paradigm into HDMapNet and get the result shown in Table III. We can observe that whether it is camera-only or camera-LiDAR fusion method, DiffMap improves the performance of HDMapNet. These illustrate our effectiveness on all types of segmentation tasks, including both long and short range detection. However, we also find that for boundaries, DiffMap does not perform well, and we conjecture that this is because the shape structure of boundaries is not fixed and has many unpredictable torsion, thereby causing difficulties in capturing the a priori structural features.
@@ IV-D Ablation Analysis
Table IV shows the impact of different downsampling factors in VQVAE on the detection results. We analyze the behavior of our DiffMap with different downsampling factors of 4, 8, 16. And We can observe that results are best when downsampling factor is set to 8x.
TABLE IV: Ablation study on the different VAE downsampling factors.
@@
Besides, we measure the effect of removing the instance-related prediction module on the model, as shown in Table V. Experiments show that adding this prediction further improves the IOU.
TABLE V: Ablation study on whether add the instance-related prediction module.
@@ IV-E Visualization
Figure 4 illustrates the comparison between DiffMap and the baseline (HDMapNet-fusion) in complex scenarios. It is evident that the segmentation results of the baseline disregard the shape properties and coherence within elements. In contrast, DiffMap demonstrates the ability to recover these problems, resulting in segmentation outputs that align well with the specifications of the map. Specifically, in cases (a), (b), (d), (e), (h), and (l), DiffMap effectively corrects inaccurately predicted pedestrian crossings. In cases (c), (d), (h), (i), (j), and (l), DiffMap completes or removes inaccurate boundaries, bringing the results closer to realistic boundary geometries. Moreover, in cases (b), (f), (g), (h), (k), and (l), DiffMap resolves the broken issue of dividers, ensuring the parallelism of neighboring elements.
@@ V CONCLUSIONS
In this paper, we introduce DiffMap, a novel approach that utilizes latent diffusion model to learn the structured priors of maps, thereby enhancing the traditional map segmentation models. Our method can serve as an auxiliary tool for any map segmentation models and its predicted results have significant improvements in both short and long range detection scenarios.
Limitation and Future work: Looking ahead, we are eager to explore the potential for direct application of diffusion models in the construction of vectorized map. Additionally, we believe that the framework is highly extensible and suitable for investigating other types of prior information, such as integrating SD map priors into the second module of DiffMap.

### v2

@@ IV-A4 Inference Details
The prediction results are obtained by performing the denoising process $20$ steps on the noisy maps under the condition of current BEV features. Considering the stochastic nature of the generative model, we perform the sampling process $3$ times, and use the mean value of sampling results as the final prediction result. We test the HDMapNet(C) with DiffMap module on one A6000 GPU, the inference FPS is 0.32.
@@ IV-B2 Instance Metrics
where $C_{1}$ and $C_{2}$ are a set of points on the predicted curve and ground-truth curve. The typical Chamfer distance(CD) is bi-directional. And the CD between the predicted curve and the ground-truth curve is given by:
@@ IV-B3 Evaluation on multiple intervals
To better explore the effectiveness of DiffMap in long-range detection, we further split the ground truth into three intervals: 0-30m, 30-60m, and 60-90m to compare with SuperFusion fairly. Subsequently, within these three intervals, we calculate the IoU, CD, and AP for different methods, aiming for a comprehensive evaluation of the detection results.
Figure 4: Qualitative results on short range map segmentation.
Figure 5: Qualitative results on long range map segmentation: Diffmap is capable of capturing the structured prior and achieves the best segmentation results. The results demonstrate the effective restoration of parallel shapes for pedestrian crossings, smoothness and continuity of dividers, and shape complementation for boundaries.
@@ IV-C Evaluation Results
To facilitate better comparison, we reproduce the predominant methods for map segmentation like HDMapNet and SuperFusion algorithms on nuScenes dataset.
Compared with Superfuion, Table I shows the comparisons of the IoU scores of semantic map segmentation (where * means the model reproduced by us). Our DiffMap shows a significant improvement on all intervals and has best results, especially in lane divider and pedestrian crossing. As shown in Table II, our method also demonstrates a significant improvement in terms of average precision (AP), verifying the effectiveness of our DiffMap. We also incorporate our DiffMap paradigm into HDMapNet and get the result shown in Table III. We can observe that whether it is camera-only or camera-LiDAR fusion method, DiffMap improves the performance of HDMapNet. They illustrate our effectiveness on all types of segmentation tasks, including both long and short range detection.
Besides we conduct experiments on cloudy, rainy and night conditions based on HDMapNet with modality C+L, the results are shown in Table IV.
However, we also find that for boundaries, DiffMap does not perform well, and we conjecture that this is because the shape structure of boundaries has many unpredictable torsion, thereby causing difficulties in capturing the a priori structural features and although there is an obvious completion correction effect on the shape, the true position is still biased.
@@ IV-D Ablation Analysis
Table V shows the impact of different downsampling factors in VQVAE on the detection results. We analyze the behavior of our DiffMap with different downsampling factors of 4, 8, 16. And We can observe that results are best when downsampling factor is set to 8x.
TABLE V: Ablation study on the different VAE downsampling factors.
@@
Besides, we measure the effect of removing the instance-related prediction module on the model, as shown in Table VI. Experiments show that adding this prediction further improves the IOU.
TABLE VI: Ablation study on whether add the instance-related prediction module.
@@ IV-E Visualization
Figure 4 illustrates the comparison between DiffMap and the baseline (HDMapNet-fusion) in complex scenarios and Figure 5 illustrates the performance of DiffMap in long-range map segmentation. It is evident that the segmentation results of the baseline disregard the shape properties and coherence within elements. In contrast, DiffMap demonstrates the ability to address these problems, resulting in segmentation outputs that align well with the specifications of the map.
@@ V CONCLUSIONS
In this paper, we introduce DiffMap, a novel approach that utilizes latent diffusion model to learn the structured priors of maps, thereby enhancing the traditional map segmentation models. Our method can serve as an auxiliary tool for any map segmentation model and its predicted results have significant improvements in both short and long range detection scenarios.
Limitation and Future work: Looking ahead, we are eager to explore the potential for direct application of diffusion models in the construction of vectorized maps. Additionally, the current model sampling latency is excessive, failing to meet real-time requirements. Therefore, further acceleration is necessary to achieve timely performance.